### v1

@@ 5.3 High-Dimensional Inverse Problem
The problem setup follows [14], which is spatially symmetric with respect to $x_{(1)}=\pi$. The source of the fluid is chosen such that ×$\nabla\times f([x_{(1)},x_{(2)}]^{T})=-\nabla\times f([2\pi-x_{(1)},x_{(2)}]^{T}).$ The observations in the inverse problem are chosen as the difference of pointwise measurements of the vorticity value $\omega([x_{(1)},x_{(2)}]^{T})-\omega([2\pi-x_{(1)},x_{(2)}]^{T})$ at $56$ equidistant points in the left domain (see Figure 6), at $T=0.25$ and $T=0.5$, corrupted with observation error $\eta\sim\mathcal{N}(0,0.1^{2}I)$. Under this set-up, both $\omega_{0}([x_{(1)},x_{(2)}]^{T})$ and $-\omega_{0}([2\pi-x_{(1)},x_{(2)}]^{T})$ will lead to the same measurements. Thus the posterior of the Bayesian inverse problem will be at least bi-modal.
@@
$\psi_{l}(x)=\begin{cases}\frac{\sin(l\cdot x)}{\sqrt{2}\pi}&l\in\mathbb{Z}_{+}^{2}\\ \frac{\cos(l\cdot x)}{\sqrt{2}\pi}&l\in\mathbb{Z}_{-}^{2}\end{cases}\quad\lambda_{l}=\frac{1}{|l|^{4}},$
and $\theta_{(l)}\sim\mathcal{N}(0,2\pi^{2})$. The KL expansion Eq. 34 can be rewritten as a sum over positive integers, where the eigenvalues $\lambda_{l}$ are in descending order. We truncate the expansion to the first $128$ terms and generate the truth vorticity field $\omega_{0}(x;\theta_{\rm ref})$ with $\theta_{\rm ref}\in\mathbb{R}^{128}$; we aim to recover the parameters based on observation data.
@@
We apply DF-GMVI with $K=3$ modes, initialized randomly according to the prior distribution and assigned equal weights. In Fig. 7, we display the true initial vorticity field, $\omega_{0}(x;\theta_{\rm ref})$, along with its mirrored field (where the mirroring of the velocity field induces antisymmetry in the vorticity field), and the three recovered initial vorticity fields, $\omega_{0}(x;m_{k})$, obtained by DF-GMVI at the 200th iteration. Modes 1 and 2 capture $\omega_{0}(x;\theta_{\rm ref})$, while mode 3 captures the mirrored field of $\omega_{0}(x;\theta_{\rm ref})$ itself. Figure 8 illustrates the relative errors in the vorticity field, optimization errors $\Phi_{R}(m_{k})$, the Frobenius norm $\lVert C_{k}\rVert_{F}$, and the Gaussian mixture weights $w_{k}$ (left to right) at each iteration. This demonstrates that DF-GMVI converges in fewer than 50 iterations. Figure 9 shows the marginal distributions of the estimated posterior for the first 16 $\theta_{(l)}$-coefficients obtained by DF-GMVI. The marginal distributions reveal clear bimodality, with the approximate posterior placing high probability mass around the true coefficients while also capturing the alternative possibility (the mirrored value). This example demonstrates DF-GMVI’s potential for effectively handling multimodal posteriors in large-scale, high-dimensional applications.
@@ 6 Conclusion
In this paper, we have explored guidelines, such as covariance positivity and affine invariance, to design stable derivative free Gaussian variational methods. Based on these guidelines, we have introduced DF-GMVI as an effective approach for solving Bayesian inverse problems. Our numerical experiments demonstrate the superior stability of DF-GMVI and its capability to approximate complex high dimensional posterior distributions. There are several promising directions for future research. On the algorithmic side, developing stable derivative free variational inference methods beyond Bayesian inverse problems Eq. 2 would be valuable. On the theoretical side, a thorough analysis of the stability and convergence properties of the natural gradient flow Eq. 20 with Gaussian mixtures and DF-GMVI for general posterior densities, including log-concave densities, could provide important insights for their practical applications.
@@ Acknowledgments
DZH acknowledges the high-performance computing platform at Peking University. YC acknowledges support from the Courant Instructorship.
Appendix A Quadrature Rule Comparison
In this section, we compare various quadrature rules, including mean-point approximation, unscented [38, 36] and cubature [2] transformations, for approximating Gaussian integrations in solving the natural gradient flow Eq. 20. We examine a 2-dimensional Gaussian mixture target density, with $\Phi_{R}(\theta)=-\log\Bigl{(}\sum_{i=1}^{3}w_{i}\mathcal{N}(\theta;m_{i},C_{i})\Bigr{)},$ where $w_{1}=0.2,\ m_{1}=\begin{bmatrix}1\\ 2\end{bmatrix},\ C_{1}=I$, $w_{2}=0.5,\ m_{2}=\begin{bmatrix}2\\ 1\end{bmatrix},\ C_{2}=I$, and $w_{3}=0.3,\ m_{3}=\begin{bmatrix}-1\\ -1\end{bmatrix},\ C_{3}=I/2$. We implement gradient and Hessian of $\Phi_{R}(\theta)$, and consider four combinations for the quadrature integrations, (1) both $\Phi_{R}$ and $\log\rho_{a}$ related terms integrated with mean-point approximation, (2) both $\Phi_{R}$ and $\log\rho_{a}$ related terms integrated with unscented transformation using weight $\sqrt{3}$ [38, 36], (3) both $\Phi_{R}$ and $\log\rho_{a}$ related terms integrated with the third-order cubature transformation [2], and (4) $\Phi_{R}$ related terms integrated with mean-point approximation and $\log\rho_{a}$ related terms with third-order cubature transformation. We use $K=10$, $\Delta t=0.02$, and solving for 1000 iterations, the convergence results are presented in Fig. 10. In the context of solving gradient flow, the mean-point approximation achieves superior convergence, even though it is not the most accurate quadrature rule. And hence the mean-point approximation is used in Definitions 3.1 and 3.2.
Figure 10: Quadrature rule comparison for the Gaussian mixture target density problem. Each panel displays the reference density and densities estimated by approximating natural gradient flow Eq. 20 with various quadrature rule combinations (from left to right). The mean $m_{k}$ (red circles) and initial mean (grey crosses) of each Gaussian component are marked. The last panel shows the total variation distance between the reference density and the densities estimated.
Appendix B Proof of Proposition 4.3
Proof B.1.
@@ Appendix A Quadrature Rule Comparison
    $=T^{-T}\sqrt{C}^{-T}(6\mathrm{Diag}(A^{T}A)+B^{T}B)\sqrt{C}^{-1}T^{-1}$
    $=T^{-T}\mathrm{\mathsf{QR}}_{\mathcal{N}}\{\nabla_{\theta}\nabla_{\theta}\Phi_{R}\}T^{-1}.$

### v3

@@ 6.3 High-Dimensional Inverse Problem
The problem setup follows [15], which is spatially symmetric with respect to $x_{(1)}=\pi$. The source of the fluid is chosen such that ×$\nabla\times f([x_{(1)},x_{(2)}]^{T})=-\nabla\times f([2\pi-x_{(1)},x_{(2)}]^{T}).$ The observations in the inverse problem are chosen as the difference of pointwise measurements of the vorticity value $\omega([x_{(1)},x_{(2)}]^{T})-\omega([2\pi-x_{(1)},x_{(2)}]^{T})$ at $35$ equidistant points in the lower-left region of the domain (see Figure 7), at $T=0.25$ and $T=0.5$, corrupted with observation error $\eta\sim\mathcal{N}(0,0.3^{2}I)$. Under this set-up, both $\omega_{0}([x_{(1)},x_{(2)}]^{T})$ and $-\omega_{0}([2\pi-x_{(1)},x_{(2)}]^{T})$ will lead to the same measurements. Thus the posterior of the Bayesian inverse problem will be at least bi-modal.
@@
$\psi_{l}(x)=\begin{cases}\frac{\sin(l\cdot x)}{\sqrt{2}\pi}&l\in\mathbb{Z}_{+}^{2}\\ \frac{\cos(l\cdot x)}{\sqrt{2}\pi}&l\in\mathbb{Z}_{-}^{2}\end{cases},\quad\lambda_{l}=\frac{1}{|l|^{4}},$
and $\theta_{(l)}\sim\mathcal{N}(0,2\pi^{2})$. The KL expansion Eq. 21 can be rewritten as a sum over positive integers, where the eigenvalues $\lambda_{l}$ are in descending order. We truncate the expansion to the first $128$ terms and generate the truth vorticity field $\omega_{0}(x;\theta_{\rm ref})$ with $\theta_{\rm ref}\in\mathbb{R}^{128}$; we aim to recover the parameters based on observation data. Therefore, the problem is ill-posed, as the number of unknown parameters exceeds the amount of observed data.
@@
We apply DF-GMVI with $K=5$ modes, initialized randomly according to the prior distribution and assigned equal weights. In Fig. 8, we display the true initial vorticity field, $\omega_{0}(x;\theta_{\rm ref})$, along with its mirrored field (where the mirroring of the velocity field induces antisymmetry in the vorticity field), and the five recovered initial vorticity fields, $\omega_{0}(x;m_{k})$, obtained by DF-GMVI at the 200th iteration. Modes 1 and 2 capture $\omega_{0}(x;\theta_{\rm ref})$, while modes 3 to 5 capture the mirrored field of $\omega_{0}(x;\theta_{\rm ref})$ itself. Figure 9 illustrates the relative errors in the vorticity field, optimization errors $\Phi_{R}(m_{k})$, the Frobenius norm $\lVert C_{k}\rVert_{F}$, and the Gaussian mixture weights $w_{k}$ (left to right) at each iteration. This demonstrates that DF-GMVI converges in fewer than 50 iterations. Figure 10 shows the marginal distributions of the estimated posterior for the first 16 $\theta_{(l)}$-coefficients obtained by DF-GMVI. The marginal distributions reveal clear bimodality, with the approximate posterior placing high probability mass around the true coefficients while also capturing the alternative possibility (the mirrored value). This example demonstrates DF-GMVI’s potential for effectively handling multimodal posteriors in large-scale, high-dimensional applications.
@@ 7 Conclusion
In this paper, we have explored guidelines, such as covariance positivity, to design stable Gaussian mixture variational inference methods. Based on these guidelines, we have introduced DF-GMVI as an effective derivative free approach for solving Bayesian inverse problems. Our numerical experiments demonstrate the superior stability of DF-GMVI and its capability to approximate complex high dimensional posterior distributions. There are several promising directions for future research. On the algorithmic side, several directions are promising for further improving practical applicability: developing stable derivative free variational inference methods beyond Bayesian inverse problems (see Eq. 2); extending the quadrature approximation defined in Definition 4.1 to higher-order schemes; and exploring adaptive strategies for determining the number of Gaussian components $K$. On the theoretical side, a thorough analysis of the stability and convergence properties of the natural gradient flow Eq. 6 with Gaussian mixtures, as well as DF-GMVI for general posterior densities, including log-concave densities, could provide insights for their practical applications.
@@ Acknowledgments
The authors acknowledge the support of the National Natural Science Foundation of China through grant 12471403, the Fundamental Research Funds for the Central Universities, and the high-performance computing platform of Peking University.
Appendix A Proof of Proposition 5.3
Proof A.1.
@@
    $=T^{-T}\sqrt{C}^{-T}(6\mathrm{Diag}(A^{T}A)+B^{T}B)\sqrt{C}^{-1}T^{-1}=T^{-T}\mathrm{\mathsf{QR}}_{\mathcal{N}}\{\nabla_{\theta}\nabla_{\theta}\Phi_{R}\}T^{-1}.$
@@
Appendix B Parameter Sensitivity Study
In this section, we study the sensitivity of DF-GMVI on hyperparameters $\alpha$ and $\Delta t$. Theoretically, the finite difference approximation in Definition 4.1 exhibits low sensitivity to relatively small $\alpha$, and Proposition 5.1 guarantees covariance positive definiteness for all $\Delta t\in(0,1)$. Empirically, we test the sensitivity by running DF-GMVI (K=40) for 200 iterations on Case E in Section 6.2, with different choices of $\alpha$ and $\Delta t$. To test the sensitivity on $\alpha$, we fix $\Delta t=0.5$ and assign $\alpha=10^{-1},10^{-3}$ and $10^{-5}$. To test the sensitivity on $\Delta t$, we fix $\alpha=10^{-3}$ and assign $\Delta t=0.25,0.50$ and $0.75$. The results are shown in Fig. 11, which demonstrate that the performance of DF-GMVI remains stable across different settings of parameters $\alpha$ (relatively small) and $\Delta t\in(0,1)$. We believe that the users can fix $\alpha=10^{-3}$ and $\Delta t=0.5$ for different Bayesian inverse problems.
Figure 11: Comparison for different choices of $\alpha$ and $\Delta t$ on the Case E of multi-dimensional problems with dimensionality 2. The left figure shows the total variation between the reference density and the densities estimated by the DF-GMVI over the iterations using different $\alpha\in\{10^{-1},10^{-3},10^{-5}\}$ with $\Delta t=0.5$, and the right figure shows the total variation using different $\Delta t\in\{0.25,0.50,0.75\}$ with $\alpha=10^{-3}$.
Appendix C Supplement Material
C.1 Natural Gradient Variational Inference
We review natural gradient variational inference from the perspective of gradient flow. Then we discuss both the Gaussian variational family and the Gaussian mixture variational family in Sections C.1.1 and C.1.2, respectively. Variational inference aims to approximate the posterior distribution by minimizing the Kullback–Leibler (KL) divergence [71, 6]
The steepest descent direction can be interpreted as
[…]
This modification leads to a different gradient flow for updating $a$ as
C.1.1 Gaussian Approximation
$\rho_{a}^{\rm G}(\theta)=\mathcal{N}(\theta;m,C)$
(33) ${\rm FIM}(a)=\begin{bmatrix}C^{-1}&\\ &X\end{bmatrix},$
(34) $XY=\frac{1}{4}C^{-1}(Y+Y^{T})C^{-1}.$
We write down the calculation of Fisher information matrix here. First we have that
$\int(\nabla_{m}\log\rho_{a}^{\rm G})^{T}(\nabla_{m}\log\rho_{a}^{\rm G})\rho_{a}^{\rm G}\mathrm{d}\theta$ $=C^{-1}\Big{(}\int(\theta-m)^{T}(\theta-m)\rho_{a}^{\rm G}\mathrm{d}\theta\Bigr{)}C^{-1}=C^{-1},$
$\int(\nabla_{m}\log\rho_{a}^{\rm G})^{T}(\nabla_{C}\log\rho_{a}^{\rm G})\rho_{a}^{\rm G}\mathrm{d}\theta$ $=0.$
For the covariance item, define a 4-th order tensor
$X=\int\big{(}\frac{1}{2}C^{-1}(\theta-m)(\theta-m)^{T}C^{-1}-\frac{1}{2}C^{-1}\bigr{)}\otimes\big{(}\frac{1}{2}C^{-1}(\theta-m)(\theta-m)^{T}C^{-1}-\frac{1}{2}C^{-1}\bigr{)}\rho_{a}^{\rm G}\mathrm{d}\theta.$
By change of variable $y=C^{-\frac{1}{2}}(\theta-m)$, we have
$X$ $=\frac{1}{4}\int\left(C^{-\frac{1}{2}}(yy^{T}-I)C^{-\frac{1}{2}}\right)\otimes\left(C^{-\frac{1}{2}}(yy^{T}-I)C^{-\frac{1}{2}}\right)\mathcal{N}(y;0,I)\mathrm{d}y,$
$X[ij,kl]$ $=\frac{1}{4}\int\sum_{p,q}C^{-\frac{1}{2}}[i,p](y_{p}y_{q}-\delta_{p,q})C^{-\frac{1}{2}}[q,j]\sum_{r,s}C^{-\frac{1}{2}}[k,r](y_{r}y_{s}-\delta_{r,s})C^{-\frac{1}{2}}[s,l]\mathcal{N}(y;0,I)\mathrm{d}y$
$=\frac{1}{4}\sum_{p,q,r,s}C^{-\frac{1}{2}}[i,p]C^{-\frac{1}{2}}[q,j]C^{-\frac{1}{2}}[k,r]C^{-\frac{1}{2}}[s,l]\int(y_{p}y_{q}-\delta_{p,q})(y_{r}y_{s}-\delta_{r,s})\mathcal{N}(y;0,I)\mathrm{d}y$
$=\frac{1}{4}\sum_{p,q,r,s}C^{-\frac{1}{2}}[i,p]C^{-\frac{1}{2}}[q,j]C^{-\frac{1}{2}}[k,r]C^{-\frac{1}{2}}[s,l](\delta_{r,p}\delta_{s,q}+\delta_{r,q}\delta_{s,p})$
$=\frac{1}{4}\left(C^{-1}[ik]C^{-1}[jl]+C^{-1}[il]C^{-1}[jk]\right).$
For all $Y\in\mathbb{R}^{N_{\theta}\times N_{\theta}}$, the act of $X$ on $Y$ is
$(XY)_{ij}$ $=\sum_{k,l}X[ij,kl]Y[kl]$
$=\frac{1}{4}\sum_{k,l}(C^{-1}[ik]C^{-1}[jl]+C^{-1}[il]C^{-1}[jk])Y[kl]$
$=\frac{1}{4}(C^{-1}YC^{-1}+C^{-1}Y^{T}C^{-1})_{ij},$
and thus $XY=\frac{1}{4}C^{-1}(Y+Y^{T})C^{-1}$.
Incorporating the Fisher information matrix Eq. 33 into Eq. 32 yields the following natural gradient flow
(35) $\dot{m}_{t}=-C_{t}\mathbb{E}_{\rho_{a_{t}}^{\rm G}}[\nabla\Phi_{R}],\qquad\dot{C}_{t}=C_{t}-C_{t}\mathbb{E}_{\rho_{a_{t}}^{\rm G}}[\nabla^{2}\Phi_{R}]C_{t}.$
(36) $\dot{C}_{t}^{-1}$ $=\mathbb{E}_{\rho_{a_{t}}^{\rm G}}[\nabla^{2}\Phi_{R}]-C_{t}^{-1}.$
The above dynamics Eq. 35 is affine invariant [14, Section 5.4.1]. Consequently, when the posterior is Gaussian, it converges exponentially fast to the posterior at a rate of $\mathcal{O}(e^{-t})$ [14, Theorem 5.6][25, 11, 9], where the exponent of the convergence rate is independent of the posterior. Furthermore, the time discretization of the natural gradient flow Eq. 35 exhibits superior stability. This stability arises from the fact that when the posterior density is log-concave, i.e., when $\nabla^{2}\Phi_{R}$ is positive semidefinite, the forward Euler discretization of Eq. 36 with $10<\Delta t<$ 1 ensures covariance positivity:
$C_{t+\Delta t}^{-1}=C_{t}^{-1}+\Delta t\bigl{(}\mathbb{E}_{\rho_{a_{t}}^{\rm G}}[\nabla^{2}\Phi_{R}]-C_{t}^{-1}\bigr{)}=(1-\Delta t)C_{t}^{-1}+\Delta t\mathbb{E}_{\rho_{a_{t}}^{\rm G}}[\nabla^{2}\Phi_{R}]\succ 0,$
provided that the approximation of $\mathbb{E}_{\rho^{\rm G}_{a_{t}}}[\nabla^{2}\Phi_{R}]$ remains positive semidefinite.
C.1.2 Gaussian Mixture Approximation
$\rho_{a}^{\rm GM}(\theta)=\sum_{k=1}^{K}w_{k}\mathcal{N}(\theta;m_{k},C_{k})$
(38) $\operatorname*{arg\,min}_{\sigma}\frac{\langle\nabla_{a}{\rm KL}[\rho_{a}^{\rm GM}\|\rho_{\rm post}],\sigma\rangle}{\sqrt{\langle\sigma,{\rm FIM}(a)\sigma\rangle}}\quad\textrm{ s.t. }\quad\sum_{k=1}^{K}\sigma_{\dot{w}_{k}}=0,$
where $\{\sigma_{\dot{w}_{k}}\}$ represents the descent directions corresponding to the weights. We introduce a Lagrangian multiplier $\lambda\in\mathbb{R}$ and write down the Karush–Kuhn–Tucker conditions of Eq. 38:
$\nabla_{d}\Big{(}\frac{\langle\nabla_{a}{\rm KL}(\rho_{a}^{\rm GM}||\rho_{\rm post}),d\rangle}{\sqrt{\langle d,{\rm FIM}(a)d\rangle}}\Bigr{)}+\lambda\nabla_{d}\Big{(}\sum_{k=1}^{K}d_{w_{k}}\Bigr{)}=0,\quad\sum_{k=1}^{K}d_{w_{k}}=0,$
where the stability condition yields
(39) $\begin{split}\langle\nabla_{a}{\rm KL}[\rho_{a}^{\rm GM}\|\rho_{\rm post}],d\rangle\,[{\rm FIM}(a)d]_{m_{k}}&=\langle d,{\rm FIM}(a)d\rangle\,\nabla_{m_{k}}{\rm KL}[\rho_{a}^{\rm GM}\|\rho_{\rm post}],\\ \langle\nabla_{a}{\rm KL}[\rho_{a}^{\rm GM}\|\rho_{\rm post}],d\rangle\,[{\rm FIM}(a)d]_{C_{k}}&=\langle d,{\rm FIM}(a)d\rangle\,\nabla_{C_{k}}{\rm KL}[\rho_{a}^{\rm GM}\|\rho_{\rm post}],\\ \langle\nabla_{a}{\rm KL}[\rho_{a}^{\rm GM}\|\rho_{\rm post}],d\rangle\,[{\rm FIM}(a)d]_{w_{k}}&=\langle d,{\rm FIM}(a)d\rangle\,\nabla_{w_{k}}{\rm KL}[\rho_{a}^{\rm GM}\|\rho_{\rm post}]-\lambda\langle d,{\rm FIM}(a)d\rangle^{\frac{3}{2}}.\end{split}$
From Eq. 39 the direction of $d$ can be determined, which leads to the following natural gradient flow:
Here $\lambda$ is determined by the constraint $\sum\dot{w}_{k}=1$. Its value depends on ${\rm FIM}(a)$.
The Fisher information matrix ${\rm FIM}(a)$ for Gaussian mixtures does not have a closed-form expression, and its inversion is computationally expensive. To improve efficiency, diagonal approximations of the Fisher information matrix have been used in the literature [15, Appendix C.8][50], leading to the following approximation:
(42) $X_{k}Y=\frac{1}{4}C_{k}^{-1}(Y+Y^{T})C_{k}^{-1}.$
We write down the derivation of Fisher information matrix approximation here. We can get Eq. 41 by only keeping the diagonal blocks of ${\rm FIM}(a)$ and approximating the diagonals under the assumptions that different Gaussian components are well separated. More precisely, for the weights items,
$\int\frac{\nabla_{w_{i}}\rho_{a}^{\rm GM}\nabla_{w_{j}}\rho_{a}^{\rm GM}}{\rho_{a}^{\rm GM}}\mathrm{d}\theta=\int\frac{\mathcal{N}_{i}\mathcal{N}_{j}}{\rho_{a}^{\rm GM}}\mathrm{d}\theta\approx\delta_{i,j}\int\frac{\mathcal{N}_{i}^{2}}{w_{i}\mathcal{N}_{i}}=\delta_{i,j}\frac{1}{w_{i}}.$
Here we substitute $\rho_{a}^{\rm GM}$ by $w_{i}\mathcal{N}_{i}$ during its integration with $\mathcal{N}_{i}$ We note that we will keep using this approximation multiple times in the following derivations. For the means items,
$\int\frac{(\nabla_{m_{i}}\rho_{a}^{\rm GM})(\nabla_{m_{j}}\rho_{a}^{\rm GM})^{\top}}{\rho_{a}^{\rm GM}}\mathrm{d}\theta$ $=\int\frac{w_{i}w_{j}\mathcal{N}_{i}\mathcal{N}_{j}C_{i}^{-1}(\theta-m_{i})(\theta-m_{j})^{\top}C_{j}^{-1}}{\rho_{a}^{\rm GM}}\mathrm{d}\theta$
$\approx\delta_{i,j}\int\frac{w_{i}^{2}\mathcal{N}_{i}^{2}C_{i}^{-1}(\theta-m_{i})(\theta-m_{i})^{\top}C_{i}^{-1}}{w_{i}\mathcal{N}_{i}}$
$=\delta_{i,j}w_{i}C_{i}^{-1}.$
For the covariances items,
$\int\frac{(\nabla_{C_{i}}\rho_{a}^{\rm GM})\otimes(\nabla_{C_{j}}\rho_{a}^{\rm GM})}{\rho_{a}^{\rm GM}}\mathrm{d}\theta$
$=$ $\int\frac{w_{i}w_{j}\mathcal{N}_{i}\mathcal{N}_{j}\big{(}C_{i}^{-1}(\theta-m_{i})(\theta-m_{i})^{\top}C_{i}^{-1}-C_{i}^{-1}\bigr{)}\otimes\big{(}C_{j}^{-1}(\theta-m_{j})(\theta-m_{j})^{\top}C_{j}^{-1}-C_{j}^{-1}\bigr{)}}{4\rho_{a}^{\rm GM}}\mathrm{d}\theta$
$\approx$ $\delta_{i,j}\int w_{i}^{2}\mathcal{N}_{i}^{2}\frac{\big{(}C_{i}^{-1}(\theta-m_{i})(\theta-m_{i})^{\top}C_{i}^{-1}-C_{i}^{-1}\bigr{)}\otimes\big{(}C_{i}^{-1}(\theta-m_{i})(\theta-m_{i})^{\top}C_{i}^{-1}-C_{i}^{-1}\bigr{)}}{4w_{i}\mathcal{N}_{i}}\mathrm{d}\theta$
$=$ $\delta_{i,j}w_{i}X_{i},$
where
$X_{i}=\frac{1}{4}\int\big{(}C_{i}^{-1}(\theta-m_{i})(\theta-m_{i})^{\top}C_{i}^{-1}-C_{i}^{-1}\bigr{)}\otimes\big{(}C_{i}^{-1}(\theta-m_{i})(\theta-m_{i})^{\top}C_{i}^{-1}-C_{i}^{-1}\bigr{)}\mathcal{N}_{i}(\theta)\mathrm{d}\theta$
is a 4-order tensor. Similar with the derivations of Eq. 33 , we have that the act of $X_{i}$ on $\forall Y\in\mathbb{R}^{N_{\theta}\times N_{\theta}}$ is given by $X_{i}Y=\frac{1}{4}C_{i}^{-1}(Y+Y^{\top})C_{i}^{-1}$.
(43) $\begin{split}\dot{m}_{k}&=-C_{k}\int\mathcal{N}_{k}(\theta)\Bigl{(}\nabla_{\theta}\log\rho_{a_{t}}^{\rm GM}+\nabla_{\theta}\Phi_{R}\Bigr{)}\mathrm{d}\theta,\\ \dot{C}_{k}&=-C_{k}\Bigl{(}\int\mathcal{N}_{k}(\theta)\bigl{(}\nabla_{\theta}\nabla_{\theta}\log\rho_{a_{t}}^{\rm GM}+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{)}\mathrm{d}\theta\Bigr{)}C_{k},\\ \dot{w}_{k}&=-w_{k}\int\Bigl{(}\mathcal{N}_{k}(\theta)-\rho_{a_{t}}^{\rm GM}\Bigr{)}\bigl{(}\log\rho_{a_{t}}^{\rm GM}+\Phi_{R}\bigr{)}\mathrm{d}\theta.\end{split}$
Here $\lambda=-\int\rho_{a}^{\rm GM}\bigl{(}\log\rho_{a_{t}}^{\rm GM}+\Phi_{R}\bigr{)}\mathrm{d}\theta$. Similar to Eq. 36, the covariance evolution equation can be rewritten as:
(44) $\frac{\mathrm{d}C_{k}^{-1}}{\mathrm{d}t}$ $=\int\mathcal{N}_{k}(\theta)\bigl{(}\nabla_{\theta}\nabla_{\theta}\log\rho_{a_{t}}^{\rm GM}+\nabla_{\theta}\nabla_{\theta}\Phi_{R}\bigr{)}\mathrm{d}\theta.$
And the weight evolution equation can be reformulated as
(45) $\frac{\mathrm{d}\log w_{k}}{\mathrm{d}t}$ $=-\int\Bigl{(}\mathcal{N}_{k}(\theta)-\rho_{a_{t}}^{\rm GM}\Bigr{)}\bigl{(}\log\rho_{a_{t}}^{\rm GM}+\Phi_{R}\bigr{)}\mathrm{d}\theta$